OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (-6, 3). @ttb123456789

OpenStudy (anonymous):

slope of perpendicular line is a negative reciprocal of the given line. then substitute the x, y coordinates of a given point and count the y-interseption then post me your answer. i need to help one anther

OpenStudy (anonymous):

i dont under stand i swear this is hurting my brain can u explain it in a more simple way pls

Nnesha (nnesha):

ok so what is your slope of this line y = 3/2x + 4 ??

OpenStudy (anonymous):

i guess

OpenStudy (anonymous):

y=-2/3x+b 3=-2/3*-6+b 3-4=b b=-1 so the answer is y=-2/3x-1

Nnesha (nnesha):

perpendicular line have opposite slope for example if your slope is a/b then for perpendicular slope flip fraction and change sign like this -b/a

OpenStudy (anonymous):

thanks for that can u help with a different question

OpenStudy (anonymous):

pls

Nnesha (nnesha):

i'll try :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but post t as new tag me ok?